OpenStudy (anonymous):

Please help! Convert to standard form:

OpenStudy (anonymous):

\[x ^{2}-6x+y ^{2}-12y+41=0\]

OpenStudy (anonymous):

i got aaaand im stuck

OpenStudy (anonymous):

the standard form of circle is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i dont know where i was going with the factoring

ganeshie8 (ganeshie8):

good try, but looks like there is a mistake

ganeshie8 (ganeshie8):

\(\large x ^{2}-6x+y ^{2}-12y+41=0\) step1 : group the variable terms separately \(\large (x ^{2}-6x)+(y ^{2}-12y) = -41\)

ganeshie8 (ganeshie8):

step2 : complete the square for the stuff inside parenthesis

ganeshie8 (ganeshie8):

6/2 = 3 so add 3^2 both sides 12/2 = 6 so add 6^2 both sides

OpenStudy (anonymous):

oh is it just (x-3)^2+(y-6)^2

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

oh and i add the 9 and 36 to -41?

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

to get 4

ganeshie8 (ganeshie8):

yess

OpenStudy (anonymous):

center is at (3,6) so i can get the h,k and put it in the standard form for a circle, but how do i find r? or is the r the 4?

ganeshie8 (ganeshie8):

again, what is the standard form of circle ?

OpenStudy (anonymous):

\[(x-h)^{2}+(y-k)^{^{2}}=r ^{2}\]

OpenStudy (anonymous):

so

ganeshie8 (ganeshie8):

and you got : \[(x-3)^2 + (y - 6)^2 = 4\]

ganeshie8 (ganeshie8):

\(\implies r^2 = 4\) \(r = ?\)

OpenStudy (anonymous):

\[(x-3)^{2}+(y-6)^{2}=r ^{2}\]

OpenStudy (anonymous):

so r=2

ganeshie8 (ganeshie8):

Correct !!

OpenStudy (anonymous):

yay, thanks!!!

ganeshie8 (ganeshie8):

yw :) good job !